OpenStudy (swaqqout_kid1):

A chemist is conducting an experiment which involves a certain chemical reaction. He begins the experiment with 866 milliliters of the chemical, which is consumed by the reaction at a varying rate that is no more than 12 milliliters per minute. Partway through the experiment, the chemist measures the amount of chemical remaining. If there are 746 milliliters of the chemical left, what is the smallest number of minutes that could have passed? 12 minutes 10 minutes 120 minutes 9 minutes @SolomonZelman

OpenStudy (swaqqout_kid1):

@BunnyBree

OpenStudy (anonymous):

You can set up a linear equation to solve for the minutes passed

OpenStudy (jhannybean):

My first guess would be to make a ratio of the information given. \[\sf \frac{866 ~\text{mL of chemical}}{12 ~\text{mL/min}}=\frac{746~\text{mL of chemical}}{x ~ \text{mins}}\]

OpenStudy (jhannybean):

Oop, i meant to write \(\sf x ~\frac{mL}{\text{mins}}\)

OpenStudy (anonymous):

\[866mm - \frac{ 12mm}{ 1 minute } \times x minutes = 746mm\]

OpenStudy (swaqqout_kid1):

?

OpenStudy (anonymous):

whoops mm...is suppose to be mL

OpenStudy (jhannybean):

xD

OpenStudy (anonymous):

haha long day sorry but both methods above work and @Jhannybean gives a better amount of significant figures.

OpenStudy (anonymous):

Solve for X in both cases